OpenStudy (anonymous):

You deposit $2,500 into a savings plan at the end of every three months. The interest rate is 5.5% compounded quarterly. Find the value of the annuity after 8 years. Do not round until the final answer. Then, round to the nearest cent.

OpenStudy (tkhunny):

P = 2500 i = 0.055 j = i/4 r = 1+j The last deposit will be just P, with no interest. The second to last deposit will be Pr, just one quarter's interest. The third to last deposit will be Pr^2, two quarter's interest. etc. All payments, accumulated together \(P + Pr + Pr^{2} + ... + Pr^{31}\). That's 32 total deposits, right? We can rephrase to \(P(1 + r + r^{2} + ... + r^{31}) = P\dfrac{1 - r^{32}}{1-r}\)

OpenStudy (anonymous):

Thank you so much for this!

OpenStudy (anonymous):

just so I can really understand this, If we rephrase it to P\[P \frac{ 1-r^{32} }{ 1-r } \] then does that mean I would just plug in \[P \frac{ 1-1.013 }{ 1-1.013 }\] ?????

OpenStudy (anonymous):

\[p \frac{ 0.54 }{ 0.013 } =41.53\]

OpenStudy (tkhunny):

?? \(r = 0.01375\) -- You chopped off WAY too many decimal places. \(r^{32} = 1.5480598641\) \(\dfrac{1-r^{32}}{1-r} = 39.8588992\) Rule of thumb: We're talking about 10s of thousands, here. Use at least 6 decimal places. \(\log_{10}(10000) = 4\). Add two (2) more for pennies, rather than dollars. Six (6) is the minimum.